OpenStudy (anonymous):

sqrt9x^5=3x^2*sqrt x ??????????

OpenStudy (mertsj):

\[\sqrt{9x^5}=3x^2\sqrt{x}\]

OpenStudy (mertsj):

Is that the problem?

OpenStudy (anonymous):

yep

OpenStudy (mertsj):

\[\sqrt{9x^4x}=3x^2\sqrt{x}\]

OpenStudy (anonymous):

its correct?

OpenStudy (anonymous):

weird

OpenStudy (mertsj):

\[3x^2\sqrt{x}=3x^2\sqrt{x}\]

OpenStudy (mertsj):

It's an identity